The speed of a boat in still water is 16 mph. The boat traveled the same distance downstream in 9 minutes that it traveled upstream in 15 minutes. Find the speed of the current....HELP!
Answer by josmiceli(6792)

c = speed of current in miles/hr
16 = speed of boat in still water (given) in miles/hr
d = distance travelled in miles
t[1] = time travelling upstream
t[2] = time travelling downstream
16 + c = speed of boat going with current in miles/hr
16 - c = speed of boat against current in miles/hr
d[1] = (rate upstream)x(time upstream)
d[2] = (rate downstream)x(time downstream)
d[1] = d[2] (given)
(16 + c)mi/hr x 9 min x 1 hr/60 min = (16 - c) x 15 min x 1 hr/60 min
notice that (mi/hr)(min)(hr/min) = mi as it should
(16 + c)(9/60) = (16 - c)(15/60)
multiply both sides by 60/3
(16 + c) x 3 = (16 - c) x 5
48 + 3c = 80 - 5c
8c = 32
c = 4 mi/hr answer
check
(16 + 4) x 3 = (16 - 4) x 5
20 x 3 = 12 x 5
60 = 60
checks